# On the extraction of Alcock-Paczyński signal from voids: a novel approach via reconstruction

### v1

@@ A.2 Sensitivity to the covariance matrices
In the analyses presented in Section 5, we used a covariance matrix estimated through jackknife resampling. This approach is consistent with standard practice in the literature and is particularly advantageous in situations where mock catalogs are unavailable.
Here, we consider an alternative method in which the covariance matrix is computed directly from the VGCF measurements obtained from the 100 mock catalogs
Figs. 10 and 11 show the normalized covariance matrices
$\text{Corr}[\boldsymbol{\xi}]_{ij}=\frac{\text{Cov}[\boldsymbol{\xi}]_{ij}}{\sqrt{\text{Cov}[\boldsymbol{\xi}]_{ii}\text{Cov}[\boldsymbol{\xi}]_{jj}}}\ .$ (29)
computed with the two methods. The 3 sub matrices along the diagonal refer to each multipoles. Cross terms are off-diagonal.
They appear qualitatively similar. The off-diagonal elements of the covariance matrix estimated from the mocks look noisier than in the jackknife case, reflecting the limited number of catalogs available (100) compared to the number of entries in the $72\times 72$ matrix (24 entries for each multipole) (Hartlap et al. 2006). Nonetheless, the diagonal elements of the two matrices are quite similar, as shown in Fig. 12.
@@
To assess the quantitative impact of the covariance matrix choice, we performed the VGCF analysis in the reconstructed space using both matrices. To correct for known biases arising from the use of a finite number of mocks, we applied the corrections proposed by Hartlap (Hartlap et al. 2006) and Percival (Percival et al. 2014) to the numerically estimated covariance matrix. The results of the two likelihood analyses are presented in the triangle plot in Fig. 13. The jackknife and mock-based covariance matrices yield consistent posterior probability distributions, resulting in unbiased estimates of the best-fit parameters.
@@
Figure 13: Posterior probability distribution of the model parameters computed with the likelihood analysis presented in Section 4, from the data of the single mock (mock 0), obtained using two different covariance matrices. Orange: jackknife resampling covariance matrix, and purple: mock covariance matrix corrected with Hartlap and Percival factors. Dark and light-shaded areas represent $68\%$ and $95\%$ confidence regions, and dashed lines indicate fiducial values of the RSD and AP parameters $\beta$ and $\varepsilon$.
@@ Appendix C Halo bias estimate
To estimate the effective linear bias values of the halos in our mock catalogs, we compute the real-space two-point auto-correlation function in each of the 100 mocks using the Landy–Szalay estimator (Landy & Szalay 1993), as implemented in the publicly available package MeasCorr (Farina et al. 2024; Guidi et al. 2023). We then compare these measurements with the BAO 2PCF model of Ross et al. (2017), using the mock-based covariance matrix, and infer the best-fit linear bias value using the BAOFitter [^5] https://gitlab.com/esarpa1/BAOFit package.

### v2

@@ A.2 Sensitivity to the covariance matrices
In the analyses presented in Section 5, we used a covariance matrix estimated through jackknife resampling. This approach is consistent with standard practice in the literature and is particularly advantageous in situations where mock catalogs are unavailable. Here, we consider an alternative method in which the covariance matrix is computed directly from the VGCF measurements obtained from the 100 mock catalogs. Figs. 10 and 11 show the normalized covariance matrices, defined as
$\text{Corr}[\boldsymbol{\xi}]_{ij}=\frac{\text{Cov}[\boldsymbol{\xi}]_{ij}}{\sqrt{\text{Cov}[\boldsymbol{\xi}]_{ii}\text{Cov}[\boldsymbol{\xi}]_{jj}}}\ ,$ (29)
computed with the two methods. The three sub matrices along the diagonal refer to each multipole. Cross terms are off-diagonal. They appear qualitatively similar. The off-diagonal elements of the covariance matrix estimated from the mocks look noisier than in the jackknife case, reflecting the limited number of catalogs available (100) compared to the number of entries in the $72\times 72$ matrix (24 entries for each multipole) (Hartlap et al. 2006). Nonetheless, the diagonal elements of the two matrices are quite similar, as shown in Fig. 12.
@@
To assess the quantitative impact of the covariance matrix choice, we performed the VGCF analysis in the reconstructed space using both matrices. To correct for known biases arising from the use of a finite number of mocks, we applied the corrections proposed by (Hartlap et al. 2006) and (Percival et al. 2014) to the numerically estimated covariance matrix. The results of the two likelihood analyses are presented in the triangle plot in Fig. 13. The jackknife and mock-based covariance matrices yield consistent posterior probability distributions, resulting in unbiased estimates of the best-fit parameters.
@@
Figure 13: Posterior probability distribution of the model parameters computed with the likelihood analysis presented in Section 4, from the data of the single mock (mock 0), obtained using two different covariance matrices. Orange: jackknife resampling covariance matrix, and purple: mock covariance matrix corrected with factors proposed by (Hartlap et al. 2006) and (Percival et al. 2014). Dark and light-shaded areas represent $68\%$ and $95\%$ confidence regions, and dashed lines indicate fiducial values of the RSD and AP parameters $\beta$ and $\varepsilon$.
@@ Appendix C Halo bias estimate
To estimate the effective linear bias values of the halos in our mock catalogs, we computed the real-space two-point auto-correlation function in each of the 100 mocks using the Landy–Szalay estimator (Landy and Szalay 1993), as implemented in the publicly available package MeasCorr (Farina et al. 2024; Guidi et al. 2023). We then compared these measurements with the BAO 2PCF model of Ross et al. (2017), using the mock-based covariance matrix, and inferred the best-fit linear bias value using the BAOFitter [^7] https://gitlab.com/esarpa1/BAOFit package.